Summary statistics for returns on two stocks X and Y are listed below.

Mean Variance
Stock X 2.83% 0.006000
Stock Y 5.98% 0.003000

The covariance of returns on stocks X and Y is 0.001500. Consider a portfolio of 80% stock X and 20% stock Y.

What is the mean of portfolio returns?

Please specify your answer in decimal terms and round your answer to the nearest thousandth (e.g., enter 12.3 percent as 0.123).
Answer #1

Ans:

We know that

when Z=aX+bY

E(Z)=aE(X)+bE(Y)

Now,

Z=0.8X+0.2Y

E(Z)=0.8*E(X)+0.2*E(Y)

=0.8*0.0283+0.2*0.0598

=0.0346

\approx 0.035